macroeconomics

Compute for the average propensity to consume and to save, and compute for the marginal propensity to consume and to save. - Also the new levels of average propensity to consume and to save after the change.

Assuming that \(Y\) or \(Y d=P 1250\) and \(C=P 1000,\) compute for APC \(\&\) APS and interpret each. Disposable income increased by \(\mathrm{P} 525\) which caused consumption spending to increase by \(\mathrm{P} 330\), compute the MPC \& MPS, and the new levels of APC \& APS after the income change. Interpret the MPC \& MPS. (20 pts)
Answer #1

APC is C/P and it shows how much of income is consumed. Here, it is \(1000 / 1250=0.8\). Thus, it implies that out of \(\$ 1\) disposable income, \(\$ .80\) is spent. Now \(\mathrm{APS}=1-\mathrm{APC}=1-0.8=0.2 .\) This shows how much of the disposable income is saved. Thus, it implies that out of \(\$ 1\) disposable income, \(\$ .20\) is saved.

MPC measures how much of increased disposable income is consumed. Here, \(\mathrm{MPC}=330 / 525=0.63 .\) This shows that when disposable income is increased by \(\$ 1,\) consumption will increase by \(\$ 0.63 .\) Now \(\mathrm{MPS}=1-\mathrm{MPC}\) and so here, it is \(1-0.63=0.37\). This implies, when disposable income is increased by \(\$ 1,\) saving is increased by \(\$ 0.37\).

New \(\mathrm{APC}=(1000+330) /(1250+525)=0.75\) and new \(\mathrm{APS}=1-0.75=0.25\)

answered by: Genersont